What is the generation time (min) of 102 E. Coli cell (CFU/g) growing logarithmically for 6 h produce 1.7 x 106 CFU/g?
1
Expert's answer
2020-07-15T18:46:03-0400

SInce, number of generation in time t is

"n=\dfrac{\ln(N)-\ln(N_o)}{\ln2}"

Thus,

"n=\dfrac{\ln(1.7\times10^6)-\ln(10^2)}{\ln2}=14"

Hence, genration time is "14/360\approx0.039" generation/min